Three businessmen A, B and C enter into partnership and agree to share the total profit in the ratio of their investment. A invested ghc450, B invested ghc150 and C invested ghc600. C received a profit of ghc24.70. Find
(i) the total profit
(ii) the profits made by A and B.
THANK YOU

Answer by Theo(13342)

A invested 450
B invested 150
C invested 600

total investment was 1200

share of A is 450/1200 = .375
share of B is 150/1200 = .125
share of C is 600/1200 = .5

C received a profit of 24.70.
that would be .5 of the total profit.
if we let x = total profit, we get .5 * x = 24.70
solve for x to get x = 24.70 / .5 = 49.40

A share of total profit = .375 * 49.4 = 18.525
B share of total profit = .125 * 49.4 = 6.175
C share of total profit .5 * 49.4 = 24.7
total profit = 49.4 as we had calculated earlier.